OpenStudy (anonymous):

The Richards Middle School has are 448 students and 32 teachers. the san miguel school has 234 students and 18 teachers. which school has a better ratio of students to teachers; that is, fewer students per teacher?

OpenStudy (anonymous):

trying to get an answer I can give to a 12 year old, to help her understand the equation and answer

OpenStudy (anonymous):

San miguel school has less students per teacher. I just divided the # of students by the # of teachers. 448/32 = 14 234/18=13

OpenStudy (anonymous):

There are 448 students for 32 teachers at Richards Middle School. There are 234 students for 18 teachers at San Miguel School. How many students are there for 1 teacher at Richards Middle School? 448/32 = 14. There are 14 students per 1 teacher at Richards Middle School. The ratio is written: 14:1 students to teachers OR as a fraction 14/1. Repeat for San Miguel. There are 14 students per 1 teacher at Richards Middle School. There are 13 students per 1 teacher at San Miguel. Which teacher has fewer students?

OpenStudy (anonymous):

thank you for your assistance!